Three charges each 5.0x10-6C are situated at he corners of an equilateral triangle of side 1.0m. Find the force that charge q3 experienced due to the other two charges.

1
Expert's answer
2022-03-21T12:47:33-0400

We know that



Force

"F_1=\frac{kq_3q_1}{a^2}"


"F_1=\frac{9\times10^9\times5\times10^{-6}\times5\times10^{-6}}{1^2}=0.225N"

"F_2=\frac{kq_2q_3}{a^2}"


"F_2=\frac{9\times10^9\times5\times10^{-6}\times5\times10^{-6}}{1^2}=0.225N"

"F_{net}=\sqrt{F_1^2+F_2^2+2F_1F_2cos60°}\\F_1=F_2=F"

We know that

"F_{net}=\sqrt{3}F"

"F_{net}=\sqrt{3}\times0.225=0.3897N"